Freon-12, CCl2F2, is used in refrigeration systems. What is the new volume (L) of a 8.0 L sample of Freon gas initially at 500 mm Hg after its pressure is changed to 2200 mm Hg at constant T? (Boyle's Law)


1
Expert's answer
2021-11-15T10:54:46-0500

"V_1 = 8.0 \;L \\

p_1 = 800 \; mm \;Hg \\

p_2 = 2200 \;mm \;Hg \\

T = constant \\

p_1V_1 = p_2V_2 \\

V_2 = \frac{p_1V_1}{p_2} = \frac{800 \times 8.0}{2200} = 2.909 \;L"

Answer: 2.909 L